# https://ku-cvlab.github.io/RAIN-GS

### v1

@@ Abstract
3D Gaussian splatting (3DGS) has recently demonstrated impressive capabilities in real-time novel view synthesis and 3D reconstruction. However, 3DGS heavily depends on the accurate initialization derived from Structure-from-Motion (SfM) methods. When trained with randomly initialized point clouds, 3DGS often fails to maintain its ability to produce high-quality images, undergoing large performance drops of 4-5 dB in PSNR in general. Through extensive analysis of SfM initialization in the frequency domain and analysis of a 1D regression task with multiple 1D Gaussians, we propose a novel optimization strategy dubbed RAIN-GS (Relaxing Accurate INitialization Constraint for 3D Gaussian Splatting) that successfully trains 3D Gaussians from randomly initialized point clouds. We show the effectiveness of our strategy through quantitative and qualitative comparisons on standard datasets, largely improving the performance in all settings.
@@
Figure 1: Effectiveness of our simple strategy. Left and right show the results from 3DGS [13] trained with dense-small-variance (DSV) random initialization (random initialization method used in the original 3DGS) and 3DGS trained with ours, respectively. Transition from 3DGS to ours simply requires sparse-large-variance (SLV) random initialization and progressive Gaussian low-pass filtering. Remarkably, each of our strategies can be implemented with a simple change in one line of code.
@@ 1 Introduction
Novel view synthesis is one of the essential tasks in computer vision and computer graphics, aiming to render novel views of a 3D scene given a set of images. It has a wide range of applications in various fields, including augmented reality and virtual reality [35], robotics [1], and data generation [8]. Neural radiance fields (NeRFs) [21] have demonstrated remarkable success in this field, learning implicit representations that capture intricate 3D geometry and specular effects solely from images. However, NeRFs’ reliance on multi-layer perceptrons (MLPs) [7, 19, 41, 27, 31] results in slow and computationally intensive volume rendering, hindering real-time applications.
@@
Recently, 3D Gaussian splatting (3DGS) [13] has emerged as a compelling alternative for both high-quality results and real-time rendering. Unlike NeRFs’ implicit representations, 3DGS models the scene using explicit 3D Gaussians. In addition, an efficient CUDA-based differentiable tile rasterization [15, 13] technique enables the rendering of the learned 3D Gaussians in real-time.
Despite its remarkable results, 3DGS exhibits a significant performance drop when trained with randomly initialized point clouds instead of those achieved from Structure-from-Motion (SfM) [13]. This limitation becomes particularly pronounced in scenarios where SfM techniques struggle to converge, such as scenes with symmetry, specular properties, and textureless regions, and limited available views. Also, due to the dependency of 3DGS on initial point clouds, SfM becomes a hard prerequisite even for situations where camera poses can be obtained through external sensors or pre-calibrated cameras [30, 9].
In this work, we start with a natural question: “Why is the initial point cloud so important in 3D Gaussian splatting?” and analyze the difference between SfM and randomly-initialized point clouds. First of all, we analyze the signals in the frequency domain on rendered images and find that SfM initialization can be interpreted as starting with a coarse approximation of the true distribution. In 3DGS, this coarse approximation acts as a foundation for subsequent refinement throughout the optimization process, preventing Gaussians from falling into local minima.
Stemming from this analysis, we conduct a toy experiment of a simplified 1D regression task, to identify the essential elements to guide the Gaussians to robustly learn the ground truth distribution from scratch. This experiment reveals that initially learning the coarse approximation (low-frequency components) of the true distribution is crucial for successful reconstruction. Similar to the SfM initialization, we find the initially learned coarse approximation acts as a guidance when learning the high-frequency components of the distribution.
Expanding upon this observation, we propose a novel optimization strategy, dubbed RAIN-GS (Relaxing Accurate INitialization Constraint for 3D Gaussian Splatting), which combines 1) a novel initialization method of starting with sparse Gaussians with large variance and 2) progressive Gaussian low-pass filtering in the rendering process. Aligned with our observation, we demonstrate that our strategy successfully guides 3D Gaussians to learn the coarse distribution first and robustly learn the remaining high-frequency components later. We show that this simple yet effective strategy largely improves the results when starting from randomly initialized point clouds without any regularization, training, or external models, effectively relaxing the requirement of accurately initialized point clouds from SfM.
@@ 2 Related work
2.0.1 Novel view synthesis.
@@
Synthesizing images from novel viewpoints of a 3D scene given images captured from multiple viewpoints is one of the essential but challenging tasks in computer vision. Neural radiance fields (NeRF) [21] have succeeded in significantly boosting the performance of this task by optimizing an MLP that can estimate the density and radiance of any continuous 3D coordinate. With the camera poses of the given images, NeRF learns the MLP by querying dense points along randomly selected rays, that outputs the density and color of each of the queried coordinates. Due to the impressive ability to understand and model complex 3D scenes, various follow-ups [3, 4, 5, 11, 17, 22, 29, 36] adopted NeRF as their baseline model and further extend the ability of NeRF to model unbounded or dynamic scenes [3, 4, 17], lower the required number of images for successful training [29, 36], learn a generalizable prior to resolve the need of per-scene optimization [5, 11], or utilize an external hash-grid to fasten the overall optimization process [22]. Although all of these works show compelling results, the volume rendering from dense points along multiple rays makes NeRF hard to apply in real-time settings achieving lower rendering rates of < 1 fps.
@@
Recently, 3D Gaussian splatting (3DGS) [13] has emerged as a promising alternative, succeeding in high-quality real-time novel-view synthesis at 1080p resolution with frame rates exceeding 90 fps. 3DGS achieves this by modeling the 3D scene with explicit 3D Gaussians, followed by an efficient and differentiable tile rasterization implemented in CUDA to accelerate the rendering process. Due to the efficient and fast architecture of 3DGS, it gained massive attention from various fields [33, 38, 20, 37], extended to model dynamic scenes [37, 20], or used as an alternative for text-to-3D tasks [33, 38] for fast generation. Nevertheless, the essential requirement of both accurate pose and initial point clouds limits the application of 3DGS to only scenes where SfM is applicable. In this work, we analyze the optimization strategy of 3DGS and present a simple yet effective change in the optimization strategy to expand the versatility of 3DGS to successfully learn the scene without accurate initialization.
2.0.2 Structure-from-Motion (SfM).
@@
SfM techniques [2, 6, 28] have been one of the most widely used algorithms to reconstruct a 3D scene. Typical SfM methods output the pose of each input image and a sparse point cloud that includes rough color and position information for each point. These methods employ feature extraction, matching steps, and bundle adjustment. For the task of novel view synthesis, NeRF [21] utilizes the estimated pose from the SfM and trains an implicit representation to model the scene. The recently proposed 3DGS [13] utilizes both the accurate pose and initial point cloud as an initialization for the position and color of 3D Gaussians, showing large performance drops when this initialization is not accessible.
@@
Despite the effectiveness of SfM in generating accurately aligned point clouds and precise camera poses, its incremental nature and the computational intensity of the bundle adjustment process significantly increase its time complexity, often to $O(n^{4})$ with respect to $n$ cameras involved [34]. Due to the high computational demand, SfM being a hard prerequisite limits the feasibility of NeRF or 3DGS for scenarios that require real-time processing. In addition, SfM often struggles to converge, such as in scenes with symmetry, specular properties, and textureless regions, and when the available views are limited.
@@ 3 Preliminary: 3D Gaussian splatting (3DGS)
Recently, 3DGS [13] has emerged as a promising alternative for the novel view synthesis task, modeling the scene with multiple 3D Gaussians [42]. Each $i$-th Gaussians $G_{i}$ represents the scene with the following attributes: a position vector $\mu_{i}\in\mathbb{R}^{3}$, an anisotropic covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$, spherical harmonic (SH) coefficients [39, 22], and an opacity logit value $\alpha_{i}\in[0,1)$. With these attributes, each Gaussian $G_{i}$ is defined in the world space [42] as follows:
@@
Since 3DGS starts learning from sparse point clouds, they utilize an adaptive density control to obtain a more accurate and denser representation. By considering if the scene is under-/over-constructed, the Gaussian undergoes a clone/split operation to represent the scene with an adequate number of Gaussians.
@@ 4 Motivation
Figure 2: Analysis of SfM initialization in 3DGS. (a) The top shows the GT image, and the bottom is the rendered image by 3DGS after only 10 steps with SfM initialization. We can observe that the rendered image is already coarsely-close to GT image. To analyze the images in the frequency domain, we randomly sample a horizontal line from the image marked in red. (b) The pixel intensity values along this line are shown, with the GT image indicated in orange and the rendered image in blue. (c) This graph visualizes the magnitude of the frequency components of (b). Since frequencies further from the middle of the x-axis represent high-frequency components, we observe that SfM provides coarse approximation of the true distribution.
4.1 SfM initialization in 3DGS
@@
To further investigate the benefits of SfM initialization, we analyze the rendered images in the frequency domain using Fourier transform [24]. In Figure 2, we randomly select a horizontal line from both the ground-truth and rendered images as shown in red in (a). The pixel intensity values along this line are visualized in (b), with the ground-truth image in orange and the rendered image in blue. These signals are then transformed into the frequency domain and presented in (c). This analysis in the frequency domain demonstrates that SfM initialization provides a coarse approximation of the underlying distribution.
As the goal of novel view synthesis is to understand the 3D distribution of the scene, it is necessary to model both low- and high-frequency components of the true distribution. To enable this, NeRF [21, 32] utilizes positional encoding to facilitate learning of high-frequency components. However, the over-fast convergence of high-frequency components hinders NeRF from exploring low-frequency components, leading NeRF to overfit to high-frequency artifcats [36]. To circumvent this problem, prior works [36, 25, 18] adopt a frequency annealing strategy, guiding NeRF to sufficiently explore the low-frequency components first.
In this perspective, starting from SfM initialization can be understood as following a similar process, as it already provides the low-frequency components (Figure 2). Expanding upon this analysis, we propose that a similar strategy that can guide 3DGS to learn the low-frequency component first is essential to enable robust training when starting with randomly initialized point clouds.
Figure 3: Toy experiment to analyze different initialization methods. This figure visualizes the result of our toy experiment predicting the target distribution using a collection of 1D Gaussians, starting from different initialization methods. Dense-small-variance (DSV) and dense-large-variance (DLV) initialize 1,000 1D Gaussians, where DLV is initialized with large variances by adding $s$ to the initial variance. Small-large-variance (SLV) initializes 15 1D Guassians with the same $s$ added to the initial variance. We can observe the over-fast convergence of high-frequency components on DSV initialization, which is resolved in the DLV initialization but fails to converge due to fluctuation. SLV initialization resolves both problems, learning the low-frequency components first and also converging to successfully model the target distribution.
4.2 Dense random initialization in 3DGS
For situations where initialized point clouds from SfM are unavailable, [13] proposes a dense-small-variance (DSV) random initialization method. They randomly sample dense point clouds within a cube three times the size of the camera’s bounding box. As the initial covariance is defined as the mean distance to the three nearest neighboring points, this leads to initializing dense 3D Gaussians with small variances.
Building upon the importance of prioritizing low-frequency learning, we conduct a toy experiment in a simplified 1D regression task to examine how DSV initialization influences the optimization process. Using 1D Gaussians provides a controlled environment isolating the effects of initialization. Specifically, we define 1D Gaussians with its mean $\mu_{i}$ and variance $\sigma_{i}^{2}$ as learnable parameters to model a random 1D signal $Y(x)$ as the ground truth distribution. By blending the $N$ Gaussians with a learnable weight $w_{i}$, we train the parameters with the L1 loss between the blended signal and $Y(x)$ as follows:
where we set $N=1,000$ for the DSV initialization setting.
As shown in the top row of Figure 3, training with the DSV random initialization exhibits a tendency towards over-fast convergence on high frequencies. This is evident in the model’s ability to capture high-frequency components after only 10 steps. However, this rapid focus on high-frequencies leads to undesired high-frequency artifacts in the final reconstruction (zoomed-in box at 1,000 steps). We hypothesize that this behavior arises from the small initial variances of the dense Gaussians, which restrict each Gaussian to model a very localized region, increasing the susceptibility to overfitting.
To verify our hypothesis, we repeat the experiment with a value $s$ added to the initial variance changing our initialization to dense-large-variance (DLV). Note that this modification effectively encourages the Gaussians to learn from wider areas. The results (middle row of Figure 3) support our hypothesis: the learned signal at 10 steps demonstrates a prioritization of low-frequency components when compared to DSV initialization. However, even with the ability to prune Gaussians by learning weights as $w_{i}=0$, the dense initialization causes fluctuations in the learned signal, preventing convergence after 1,000 steps. These observations highlight a crucial point: while learning from wider areas (enabled by a large variance) is necessary to prioritize low-frequency components, a dense initialization can still lead to instability and convergence issues. More detailed settings and results of the toy experiment are provided in Section C in the supplementatry materials.

### v2

# Relaxing Accurate Initialization Constraint for 3D Gaussian Splatting

@@ Abstract
3D Gaussian splatting (3DGS) has recently demonstrated impressive capabilities in real-time novel view synthesis and 3D reconstruction. However, 3DGS heavily depends on the accurate initialization derived from Structure-from-Motion (SfM) methods. When the quality of the initial point cloud deteriorates, such as in the presence of noise or when using randomly initialized point cloud, 3DGS often undergoes large performance drops. To address this limitation, we propose a novel optimization strategy dubbed RAIN-GS (Relaxing Accurate INitialization Constraint for 3D Gaussian Splatting). Our approach is based on an in-depth analysis of the original 3DGS optimization scheme and the analysis of the SfM initialization in the frequency domain. Leveraging simple modifications based on our analyses, RAIN-GS successfully trains 3D Gaussians from sub-optimal point cloud (e.g., randomly initialized point cloud), effectively relaxing the need for accurate initialization. We demonstrate the efficacy of our strategy through quantitative and qualitative comparisons on multiple datasets, where RAIN-GS trained with random point cloud achieves performance on-par with or even better than 3DGS trained with accurate SfM point cloud.
@@
Figure 1: Effectiveness of our simple strategy. Left and right show the results from 3DGS [1] and ours trained with randomly initialized point cloud respectively. Transition from 3DGS to ours simply requires our strategy consisted of sparse-large-variance (SLV) random initialization, progressive Gaussian low-pass filtering, and adaptive bound-expanding split (ABE-Split) algorithm.
@@ 1 Introduction
Novel view synthesis is one of the essential tasks in computer vision and computer graphics, aiming to render novel views of a 3D scene given a set of images. It has a wide range of applications in various fields, including augmented reality and virtual reality [2], robotics [3], and data generation [4]. Neural radiance fields (NeRFs) [5] have demonstrated remarkable success in this task, learning implicit representations that capture intricate 3D geometry and specular effects solely from images. However, NeRFs’ reliance on multi-layer perceptrons (MLPs) [6, 7, 8, 9, 10] results in slow and computationally intensive volume rendering, hindering real-time applications.
@@
Recently, 3D Gaussian splatting (3DGS) [1] has emerged as a compelling alternative for both high-quality results and real-time rendering. Unlike NeRFs training an MLP to model the scene with an implicit representation, 3DGS models the scene using explicit 3D Gaussians. The learned 3D Gaussians are then projected to the image space using an efficient CUDA-based differentiable tile rasterization [11, 1] technique which enables the rendering of the Gaussians in real-time.
Despite its remarkable results, 3DGS requires an additional input of initial point cloud compared to NeRFs. The quality of the initial point cloud is one of the essential requirements of 3DGS, showing large performance drops when trained with randomly initialized point cloud [1]. A similar type of performance degradation can occur in real-world scenarios, specifically in scenes where SfM techniques struggle to converge, such as scenes with symmetry, and textureless regions [12, 13]. Also, accurate initial point cloud become a strict requirement even for situations where camera poses can be obtained through external sensors [14, 15] or pre-defined such as in text-to-3D generation [16, 17]. In these cases, where the initial point cloud from SfM may contain noise or even unavailable, 3DGS requires an additional module to obtain the point cloud, adding extra computation cost [18, 19].
In this work, we start with a natural question: “Why is accurate initial point cloud so important in 3D Gaussian splatting?” and conduct an in-depth analysis of the original 3DGS optimization scheme and SfM initialization. Our analyses reveal that the original 3DGS optimization scheme struggles to transport Gaussians further from their initialized locations, leading to performance degradation when starting from random or noisy SfM initialization. In addition, the analysis of the SfM initialization in the frequency domain reveals that SfM initialization provides the low-frequency components of the true distribution. From the provided low-frequency components, 3DGS clones and splits existing Gaussians, which is then used to learn the remaining high-frequency details of the scene [1]. In this process, the low-frequency components guide the Gaussians to successfully learn the target distribution in a coarse-to-fine manner utilizing the clone/split process, without the need to transport the Gaussians to correct locations. Based on this observation, we propose that to train 3DGS with noisy or random point cloud, we need an effective strategy that can transport the Gaussians further from their initial positions and encourage the Gaussians to prioritize the learning of the coarse structure information, which will then guide the overall optimization process to learn in a similar coarse-to-fine manner.
Based on our analysis of accurate initial point cloud, we propose a novel optimization strategy called RAIN-GS (Relaxing Accurate INitialization Constraint for 3D Gaussian Splatting), which is composed of three key components: 1) a novel initialization method starting with sparse Gaussians with large variance, 2) progressive Gaussian low-pass filtering utilized in the rendering process, and 3) a novel adaptive bound-expanding split (ABE-Split) algorithm utilized in the adaptive density control. We demonstrate that RAIN-GS effectively relaxes the requirement of accurate initialization for 3DGS, as applying RAIN-GS to random point cloud successfully priorities the learning of low-frequency components and guides the Gaussians to learn in a coarse-to-fine manner, and enables the Gaussians to traverse further from their initialized locations. After optimization, RAIN-GS applied to random point cloud achieves on-par or even better results compared to 3DGS trained with SfM. By effectively relaxing the requirement of initial point cloud, RAIN-GS opens up new possibilities for applying 3DGS to settings where obtaining accurate initial point cloud is challenging.
@@ 2 Related work
2.1 Novel view synthesis
@@
Neural radiance fields (NeRF) [5] have succeeded in significantly boosting the performance of novel view synthesis by optimizing an MLP that can estimate the density and radiance of any continuous 3D coordinate. With the camera poses of the given images, NeRF learns the MLP by querying dense points along randomly selected rays, that outputs the density and color of each of the queried coordinates. Various follow-ups [20, 21, 22, 23, 24, 25, 26, 27] adopted NeRF as their baseline model and further extend the ability of NeRF to model unbounded or dynamic scenes [20, 21, 24], lower the required number of images for successful training [26, 27], learn a generalizable prior to resolve the need of per-scene optimization [22, 23], or utilize an external hash-grid to fasten the overall optimization process [25]. Although all of these works show compelling results, the volume rendering from dense points along multiple rays makes NeRF hard to apply in real-time settings achieving lower rendering rates of < 1 fps.
@@
Recently, 3D Gaussian splatting (3DGS) [1] has been recognized as a new solution for the task of novel view synthesis, succeeding in high-quality real-time novel-view synthesis at 1080p resolution with frame rates exceeding 90 fps. 3DGS achieves this by modeling the 3D scene with explicit 3D Gaussians, followed by an efficient and differentiable tile rasterization implemented in CUDA to accelerate the rendering process. Due to the efficient and fast architecture of 3DGS, it gained massive attention from various fields [28, 17, 29, 16], extended to model dynamic scenes [29, 28], or used as an alternative for text-to-3D tasks [16, 17] for fast generation. Nevertheless, the essential requirement of both accurate pose and initial point cloud limits the application of 3DGS to only scenes where SfM can provide accurate point cloud. In this work, we analyze the optimization strategy of 3DGS and present a simple yet effective change in the optimization strategy to expand the versatility of 3DGS to successfully learn the scene without accurate initialization.
2.2 Structure-from-Motion (SfM)
@@
SfM techniques [30, 31, 32] have been one of the most widely used algorithms to reconstruct a 3D scene. Typical SfM methods output the pose of each input image and a sparse point cloud that includes rough color and position information for each point. These methods employ feature extraction, matching steps, and bundle adjustment. For the task of novel view synthesis, NeRF [5] utilizes the estimated pose from the SfM and trains an implicit representation to model the scene. The recently proposed 3DGS [1] utilizes both the accurate pose and initial point cloud as an initialization for the position and color of 3D Gaussians, showing large performance drops when the initial point cloud becomes noisy or not accessible.
@@
Despite the effectiveness of SfM in generating accurately aligned point cloud and precise camera poses, its incremental nature and the computational intensity of the bundle adjustment process significantly increase its time complexity, often to $O(n^{4})$ with respect to $n$ cameras involved [33]. Due to the high computational demand, SfM being a hard prerequisite limits the feasibility of NeRF or 3DGS for scenarios that require real-time processing. In addition, SfM often struggles to converge, such as in scenes with symmetry, specular properties, and textureless regions, and when the available views are limited. In these cases, an additional module is required to provide accurate initial point cloud to 3DGS, which can take longer times than SfM [19].
@@ 3 Preliminary: 3D Gaussian splatting (3DGS)
Recently, 3DGS [1] has emerged as a promising alternative for the novel view synthesis task, modeling the scene with multiple 3D Gaussians [34]. Each $i$-th Gaussian $G_{i}$ represents the scene with the following attributes: a position vector $\mu_{i}\in\mathbb{R}^{3}$, an anisotropic covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$, spherical harmonic (SH) coefficients [35, 25], and an opacity logit value $\alpha_{i}\in[0,1)$. With these attributes, each Gaussian $G_{i}$ is defined in the world space [34] as follows:
@@
Since 3DGS starts learning from sparse point cloud, they utilize an adaptive density control to obtain a more accurate and denser representation. By considering if the scene is under-/over-constructed, each of the Gaussians undergoes a clone/split operation. Gaussians that have small covariance are cloned and Gaussians considered to be too large are splitted into two new smaller Gaussians, where the means of new Gaussians are sampled from the probability density function (PDF) of the original Gaussian. When the Gaussian is considered to be too small, the Gaussian is cloned and simply located in the mean of the original Gaussian.
@@ 4 Motivation
Although point cloud can be noisy or unavailable in real-world scenarios [12, 13], 3DGS shows large performance drops depending on the accuracy of the initial point cloud [1]. To understand the large performance gap of 3DGS, we conduct an in-depth analysis of the original 3DGS optimization scheme and the analysis of SfM initialization in the frequency domain. Our analyses reveal two important characteristics of the optimization scheme of 3DGS: 1) the optimization scheme of 3DGS struggles to transport Gaussians from their initialized locations and 2) the coarse structure information (low-frequency components) provided by the accurate initialization enables the adaptive density control method of 3DGS to robustly model the remaining fine details of the scene in a coarse-to-fine manner.
3DGS lacks the ability to transport Gaussians.
To represent and learn the scene with explicit 3D Gaussians, 3DGS first initializes the Gaussians $G_{i}$ in the world space, whose means $\mu_{i}$ are defined by the initial point cloud. The point cloud can be either achieved from SfM or initialized randomly.
As mentioned in [36], the process of fitting a 3DGS model is similar to fitting a Gaussian Mixture Model (GMM), which is well-known for being non-convex and generally solved with the Expectation-Maximization (EM) algorithm [37]. They further note that, similar to the EM algorithm, training 3DGS from randomly initialized point cloud becomes prone to falling into local minima due to two main reasons. 1) The Gaussians can only receive gradients close to their means, mostly from the range not exceeding the distance of a few standard deviations, and 2) there is no existing path for the Gaussians that will decrease the loss monotonically.
Although [36] only analyzes the case of starting from random initialization, we verify that Gaussians can also easily fall into local minima when SfM-initialized point cloud become noisy. As shown in Table 2, we find that adding a small constant noise or adding a small noise $\epsilon$ sampled from a normal distribution ($\epsilon\sim\mathcal{N}(0,1))$ to the SfM-initialized point cloud, leads to large performance drops. Based on these observations, we hypothesize that the optimization scheme of 3DGS lacks the ability to correct or move the positions of the Gaussians. We empirically verify our hypothesis by calculating the average distance each Gaussian traversed after optimization, as shown in Table 2. It can be seen that the average distance each Gaussian moved is close to zero, indicating that the optimization scheme of 3DGS lacks the ability to move Gaussians, which can lead to the failure of capturing objects located far from the initial positions of the Gaussians. This emphasizes the need for a strategy that can enable the Gaussians to transport further from their initialized locations, in order to successfully train 3DGS from sub-optimal initializations.
Table 1: Effect of noise on the initial SfM point cloud.
@@
SfM 27.205 0.815 0.214
SfM + ${\epsilon}$ 24.944 0.771 0.270
SfM + constant 26.204 0.761 0.277
Table 2: Analysis of Gaussians movement on Truck scene.
3DGS (SfM) 3DGS (Random) Ours (Random)
Means 0.349 0.184 12.100
Stds 0.726 0.297 18.600
Top 1% 3.418 1.524 82.721
Accurate initialization guides 3DGS to learn in a coarse-to-fine manner.
Figure 2: Analysis of SfM initialization in 3DGS. (a) The top shows the GT image, and the bottom is the rendered image by 3DGS after only 10 steps with SfM initialization. We can observe that the rendered image is already coarsely-close to GT image. We randomly sample a horizontal line from the image marked in red. (b) The pixel intensity along this line are shown, with the GT indicated in blue and the rendered image in orange. (c) This graph visualizes the magnitude of the frequency components of (b). Since frequencies further from the middle of the x-axis represent high-frequency components, we observe that SfM provides coarse approximation of the true distribution.
@@
To further investigate the benefits of SfM initialization, we analyze the rendered images in the frequency domain using Fourier transform [38]. As shown in Figure 2, the analysis in the frequency domain demonstrates that SfM initialization provides a coarse approximation of the target distribution.
As the goal of novel view synthesis is to understand the 3D distribution of the scene, it is necessary to model both low- and high-frequency components of the true distribution. However, prior NeRF frameworks [39, 40, 27] argue that NeRF is prone to overfitting and naïve optimization leads to over-fast convergence of high-frequency components, expressed with high-frequency artifacts in the rendered image. To circumvent this problem, they adopt a coarse-to-fine learning strategy, which regularizes NeRF to learn the low-frequency components first. Similarly, prior works [41, 42] utilizing GMMs for the task of point cloud registration or generation also mention that naïve fitting of GMMs can result in converging to local minima. In order to robustly train GMMs, they also adopt a coarse-to-fine strategy, implemented by starting with a small number of Gaussians and recursively increasing the number of total Gaussians. In both NeRFs and GMMs, coarse-to-fine strategy guides the network to learn more robustly, leading to better performance.
In this perspective, starting the optimization of 3DGS from SfM-initialized point cloud can be understood as benefitting from a similar coarse-to-fine process, where SfM provides the low-frequency components (Figure 2), and the adaptive density control method of 3DGS adds the Gaussians to learn the remaining high-frequency details. Based on our observations, the success of 3DGS from accurate initialization can be attributed to the low-frequency components guiding the overall training process, preventing the Gaussians from falling into local minima. This highlights the need for a strategy that can prioritize the learning of the low-frequency components even from sub-optimal initializations, which will then be used to guide the remaining optimization process of 3DGS.